4. Problems and Applications Q4 An economy consists of three workers: Carlos, Felix, and Larry. Each works 10 hours a day and can produce two services: mowing lawns and washing cars. In an hour, Carlos can either mow 2 lawns or wash 1 car; Felix can either mow 1 lawn or wash 1 car; and Larry can either mow 1 lawn or wash2 cars. For each of the scenarios listed in the following table, determine how many lawns will be mowed and how many cars will be washed per day and enter these values into the corresponding row Scenario All three spend all their time mowing lawns. (A) All three spend all their time washing cars. (B) All three spend half their time on each activity. (C) Carlos spends half his time on each activity, while Felix only washes cars and Larry only mows lawns. (D) Lawns Mowed Cars Washed In the following table, identify the opportunity cost of washing cars for each worker. Worker Carlos Felix Larry Opportunity Cost of Washing Cars Assume that the resources best suited to producing a particular service are preferentially used in the production of that service and that as the economy moves down along the production possibilities frontier, one worker at a time is transferred from mowing lawns to washing cars. Using the blue points (circle symbol), graph the production possibilities frontier (PPF) for this economy on the following graph. Then use the black point (plus symbol) to identify point A, the green point (triangle symbol) to identify point B, the orange point (square symbol) to identify point C, and the purple point (diamond symbol) to identify point D on the graph.

Assume that the resources best suited to producing a particular service are preferentially used in the production of that service and that as the economy moves down along the production possibilities frontier, one worker at a time is transferred from mowing lawns to washing cars. Using the blue points (circle symbol), graph the production possibilities frontier (PPF) for this economy on the following graph. Then use the black point (plus symbol) to identify point A, the green point (triangle symbol) to identify point B, the orange point (square symbol) to identify point C, and the purple point (diamond symbol) to identify point D on the graph. 50 T 45 PPF 40 8 35 Σ30 25 ž 20 О 15 10 0 5 10 15 20 25 30 35 40 45 50 Quantity of Cars Washed True or False: The production possibilities frontier consists of straight-line segments, rather than being smoothly bowed outward throughout, because each worker faces a constant trade-off between mowing lawns and washing cars O True False

Indicate whether each of the following allocations is efficient or inefficient. Allocation Efficient Inefficient
Answer #1
Lawns per hour Car wash per hour
Carlos 2 1
Felix 1 1
Larry 1 2

Total hours they can work individually = 10 hours

PPF 45 2 20 15 0 5 10 15 20 25 3035 4045 Cars Washed

Scenario Lawns Car
All three spend their time mowing lawn A 20+10+10=40 0
All three spend their time washing cars B 0 10+10+20 = 40
All three spend half their time in each activity C 10+5+5 = 20 5+5+10 = 20
Carlos spends half his time on each activity, Felix only washes car and Larry only mows lawn D 10+0+10=20 5+10+0=15

Opportunity cost

Opportunity Cost
Worker opportunity cost of washing cars
Carlos 2 Lawns
Felix 1 Lawn
Larry 0.5 Lawn

True. Refer the picture attached. We have a straight line.

Efficient Inefficient
A Efficient
B Efficient
C Efficient
D Inefficient

D is inefficient since it is below the PPF.

Please contact if having any query thank you.